Janet invested $31,000, part at 10% and part at 1%. If the total interest at the end of the year is $1,390, how much did she invest at 10%?
Answer by checkley75(3666)

.10X+.01(31,000-X)=1.390
.10X+310-.01X=1,390
.09X=1,390-310
.09X=1,080
X=1,080/.09
X=12,000 INVESTED @ 10%.
PROOF
.10*12,000+.01*19,000=1,390
1200+190=1,390
1,390=1,390